OpenStudy (anonymous):

2a-6 divided by 3a-9

OpenStudy (anonymous):

Note that \(\large 2a-6=2(a-3)\) and \(\large 3a-9 = 3(a-3)\). So \(\large \dfrac{2a-6}{3a-9}=\ldots\) Can you take things from here? :-)

OpenStudy (nirmalnema):

2a-6 / 3a-9 2(a-3)/3(a-3) =2/3